# Counterfactual Visual Explanation via Causally-Guided Adversarial Steering

### v1

@@ 4.2 Causally-Guided Post Refinement
$\arg\min_{x^{\prime\prime}}L(F(x^{\prime\prime});y^{\prime})+\lambda||x^{\prime\prime}-x||_{1},$ (7)
@@
The mask indicates which pixels need to be modified through $l_{1}$ loss and which do not. Intuitively, we need to combine $x^{\prime}$ generated in the last subsection and the original image $x$ with the mask to achieve causal perturbation in CEs with well-preserved pixels in the rest parts. However, directly combining $x^{\prime}$ and $x$ using the mask is not feasible, as it would result in clearly visible region boundaries in the CE. In our work, we leverage the image inpainting technique [68] to tackle this problem. The inpainting process is shown as follows:
@@ 5.1 Experimental Setup
Datasets We conducted extensive experiments on 8 datasets in total. We first evaluate our framework on the CelebA dataset [69], using images of size 128$\times$128 and focusing on two attributes which constitute two sub-datasets: “Smile” and “Age”. The CelebA-HQ [70] dataset is used with the same attributes and is also constructed into two sub-datasets, but with images of size 256$\times$256. The other datasets include the traffic scene dataset BDD-OIA [71], with images of size 512$\times$256, and three sub-datasets extracted from ImageNet [72], each containing two classes. These three sub-datasets are: Cougar vs. Cheetah, Sorrel vs. Zebra, and Egyptian Cat vs. Persian Cat. All experiments were conducted on NVIDIA RTX 4090 and RTX A6000 GPUs.
Baselines We selected four recent and state-of-the-art methods as our baselines: STEEX [8], DiVE [10], DiME [9], and ACE [11]. Among them, DiVE includes an extended variant called DiVE100, in which the number of optimization steps is reduced to 100.
Models Weights and Architectures To ensure a fair comparison, we use the same pre-trained DDPM weights as those used in the baselines for each dataset. For ImageNet, we use the ResNet-50 model pre-trained by PyTorch [73] as the classifier. For the other datasets, we use DenseNet-121 as the classifier and adopt the pre-trained weights provided by the baselines. Additional training details, including hyperparameter settings, are provided in Appendix A.
Evaluation Metrics We evaluate our method and the baselines across multiple dimensions, each assessed using several metrics. The first evaluation dimension is validity, which we measure using Flip Rate (FR) and Counterfactual Transition (COUT). FR denotes the proportion of samples for which the counterfactual successfully causes the classifier to change its prediction. COUT measures the transition scores between the original and counterfactual images, with values ranging from $-1$ to $1$. The second evaluation dimension focuses on sparsity and proximity, which we assess using Learned Perceptual Image Patch Similarity (LPIPS) [74], Correlation Difference (CD) [9], and SimSiam Similarity (S3) [75]. LPIPS measures the perceptual similarity between two images based on human visual perception. CD computes the average number of non-target attributes that are unintentionally altered across all samples. S3 evaluates the cosine similarity between the original and counterfactual images using SimSiam as the encoding network. The third evaluation dimension is realism, which we measure using Fréchet Inception Distance (FID) [76] and split FID (sFID) [11]. Since counterfactual images typically share most pixels with the original images, FID tends to be underestimated and may fail to reflect their true differences. To address this issue, sFID is proposed. It splits both the original and counterfactual image sets into two subsets, computes FID in a cross-wise manner, and repeats this process ten times to obtain an average score. All results reported in this paper are averaged over multiple experimental runs. We omit standard deviations as they are negligible compared to the corresponding mean values.
Table 1: Evaluation Results on CelebA: Smile.
@@
DiVE - - - - - 29.4 -
DiVE100 - - - - 2.34 36.8 -
STEEX - - - - - 10.2 -
DiME 97.60% -0.0451 0.0896 0.9158 1.99 12.3 25.48
ACE 99.05% 0.6547 0.0273 0.9822 2.05 4.2 26.38
Ours 99.40% 0.7325 0.0138 0.9834 2.09 3.8 25.70
Table 2: Evaluation Results on CelebA: Age
@@
DiVE - - - - - 33.8 -
DiVE100 - - - - - 39.9 -
STEEX - - - - - 11.8 -
DiME 73.10% 0.0381 0.0938 0.9158 3.81 20.2 34.37
ACE 99.00% 0.5441 0.0326 0.9843 3.37 4.3 26.47
Ours 99.95% 0.6908 0.0167 0.9830 2.98 4.6 26.26
Table 3: Evaluation results on BDD-OIA.
Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ FID $\downarrow$ sFID $\downarrow$
DiME 60.50% -0.4865 0.1119 0.9638 62.30 225.17
ACE 99.90% 0.7278 0.0318 0.9956 9.50 91.70
Ours 99.90% 0.8240 0.0150 0.9952 13.85 92.48
@@
Datasets Methods FR $\uparrow$ COUT $\uparrow$ LIPIPS $\downarrow$ S3 $\uparrow$ FID $\downarrow$ sFID $\downarrow$
Sorrel- Zebra ACE 76.67% 0.0967 0.1670 0.9285 77.76 124.66
Ours 90.00% 0.5442 0.1714 0.9367 74.58 114.66
Cougar- Cheetah ACE 76.67% 0.2059 0.1809 0.9625 108.67 147.79
Ours 93.33% 0.3390 0.1855 0.9693 105.49 146.04
Persian Cat- Egyptian Cat ACE 80.00% 0.2326 0.1531 0.8939 139.65 227.07
Ours 83.33% 0.3424 0.1543 0.8759 134.57 213.21
@@ 5.2 Quantitative Analysis
In this section, we conduct a quantitative analysis of all datasets with results summarized in Table 1 through Table 6. We highlight the best performance in bold and the second-best in italics. The results for the baselines DiVE, DiVE100, and STEEX, we extract the results from the original papers. For the data not reported in the original papers, we leave the corresponding entries blank in the tables.
Tables 1 and  2 show the results on the CelebA dataset. As observed, our method achieves the best performance on 5 out of the 7 metrics and ranks second on the remaining two. More specifically, our method achieves the highest FR and COUT scores, while also attaining the lowest LPIPS and S3 values, along with relatively low FID and sFID scores. This indicates that our method can achieve the highest flip rate with minimal modifications. In other words, it attains the highest validity while preserving both sparsity and proximity. Due to the similarity between the CelebA-HQ and CelebA datasets, we include the results for CelebA-HQ in Appendix B. Table 3 presents the results on the BDD-OIA dataset, which is the most complex among all, containing rich traffic-related elements. Due to space constraints, we omit the results of DiVE and STEEX on this dataset with a similar observation in other benchmarks. Our method did not improve the FR on this dataset because the FR is already near its upper bound. However, it still achieves the highest COUT score without incurring significant degradation in FID and sFID. This suggests that CECAS is capable of identifying more effective counterfactual transition directions, leading to increased classifier confidence in the target label prediction. Table 6 reports the results on the three ImageNet subsets. Due to the page limit, we focus on the comparison between our framework and the most competitive baseline ACE. Apart from marginally lower performance in LPIPS and S3 on the final subset, our method significantly outperforms ACE on all other metrics.
Figure 3: Qualitative results on CelebA dataset. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
Figure 4: Qualitative results on BDD-OIA dataset. Green circles highlight the counterfactual changes. The right column for each transition direction shows a zoomed-in version of the left column.
Figure 5: Qualitative results on ImageNet dataset. Green circles highlight the counterfactual changes.
@@ 5.3 Qualitative Analysis
In this section, we perform case studies to straightforwardly evaluate the quality of the generated counterfactual images from a human perceptual perspective. Due to the complexity of the BDD-OIA and ImageNet datasets, it is difficult to discern the differences between counterfactual images generated by our method and those produced by other methods through visual inspection. Therefore, for these two datasets, we only present the results of our method. For the other datasets, we compare the counterfactual images generated by our approach with those produced by the best-performing method, ACE, to demonstrate the visual superiority of our method. As shown in Figure 3, ACE introduces inexplicable color and shape distortions, as well as unintended attributes in the generated images. For example, in the transformation from “young” to “old,” some female samples are incorrectly rendered with dark patches on the chin—an artifact caused by the spurious correlation between beard and aging in male samples. This results in unmeaningful CEs. In contrast, our method effectively mitigates such artifacts by avoiding reliance on spurious cues. Figures 4 and 5 present the results on complicated datasets BDD-OIA and ImageNet, respectively. The green circles in the images highlight the necessary and human-perceptible changes made to achieve counterfactual visual explanations. We include the results for CelebA-HQ in Appendix B for the same reason. More detailed results for the other datasets are also provided in Appendix B.
@@ 5.4 VLM Evaluations
Vision-Language Models (VLMs), through large-scale pretraining on image–text pairs, demonstrate strong reasoning capabilities in tasks such as visual question answering and image captioning [77]. Overcoming the limitation of traditional CE metrics that only provide shallow numerical evaluations and fail to capture the semantic meaning of images, VLMs provide more human-like evaluations by leveraging their rich prior knowledge in pertaining [78]. To better suit the specific nature of counterfactual image generation, we carefully design unique prompts with evaluation criteria covering validity, sparsity, proximity, and realism. The prompt design for each dataset can be found in Appendix C. Notably, sparsity in our prompt is partially assessed from a causal perspective, examining whether the modified regions in the image are necessary or causally relevant, while ensuring that unrelated parts remain unchanged. In the context of our task, we utilize GPT-4o to assign scores ranging from 1 to 10 for each of the four evaluation dimensions. Given that these dimensions hold equal importance in counterfactual generation, the final score is computed as the average across all four. As shown in Table 6, our method outperforms the baselines in 7 out of the 8 datasets. Combined with the powerful visual and textual reasoning capabilities of VLMs, we have strong reason to believe that the counterfactual images generated by our method offer the highest level of interpretability.
@@ 5.5 Parameters Study
Figure 6: Line plots of FR and FID with respect to varying values of $\alpha$.
In this section, we analyze the impact of the hyperparameter $\alpha$ on counterfactual image generation. We present results on the CelebA: Smile dataset, while results on the other datasets are provided in Appendix D. As shown in Figure 6, FID initially decreases and then increases as $\alpha$ increases. This is because, in the early stages of increasing $\alpha$, the constraint on changes to spurious factors becomes stronger, reducing unintended modifications and thereby making the generated counterfactual images closer to real ones. However, as $\alpha$ continues to increase, the counterfactual generation process begins to overemphasize the alignment of spurious distributions. As a result, regions that originally required slight modifications become overly constrained, making the generated images more unreal. Consequently, the FID score gradually increases again. Meanwhile, the increase in $\alpha$ leads to a decline in FR, this is reasonable since when $\alpha=0$, the flip rate is highest as the optimization focuses solely on flipping the prediction. However, this comes at the cost of poor counterfactual quality, often producing unrealistic images. Therefore, it is crucial to strike a balance between flip effectiveness and visual realism. Fortunately, with a properly chosen $\alpha$, our method effectively achieves this balance, producing counterfactuals with both high flip rates and high image quality.
@@ 6 Conclusion
In this paper, we propose a novel framework, CECAS, which innovatively designs a causally-guided adversarial method to generate counterfactual visual explanations that constrain less informative changes on spurious factors. The framework can be applied to any explained DNN models. Experimental results demonstrate that our method consistently outperforms existing approaches in quantitative evaluation, qualitative analysis, and VLM-based assessment. CECAS achieves a better balance among the four key properties of counterfactual explanations: validity, sparsity, proximity, and realism. Therefore, we argue that the counterfactual explanations generated by our method offer a more causally grounded understanding of the classifier’s decision-making and prediction mechanisms.
Limitations While our method focuses on factors that are causally related to the label, it does not explicitly model more fine-grained causal relationships between more variables, which could further enhance consistency in counterfactual generation. Exploring such nuanced causal structures remains a valuable direction for future work. Additionally, developing techniques to automatically balance competing objectives would be important toward more adaptive and generalizable CE methods.
@@
A. Training Details
Data Preparation The eight datasets discussed in the main text are categorized into two types: binary datasets and multi-label datasets. CelebA Smile, CelebA Age, CelebA-HQ Smile, CelebA-HQ Age, and BDD-OIA belong to the first category, while Cougar vs. Cheetah, Sorrel vs. Zebra, and Egyptian Cat vs. Persian Cat fall into the latter. For binary-label datasets, if the original predicted label is $y$ ($y$ can either be 0 or 1), then the target label is set to $1-y$. For multi-label datasets, since ImageNet contains 1,000 classes, the prediction and target labels differ for each dataset subset. Specifically, Cougar vs. Cheetah: predicted label is 286, corresponding to “cougar”; target label is 293, corresponding to "cheetah". We adopt this setting to perform a transition from the “cougar” class to the “cheetah” class. Sorrel vs. Zebra: The predicted label is 339, which corresponds to "sorrel"; the target label is 340, which corresponds to "zebra". We adopt this setting to perform a transition from the “sorrel” class to the “zebra” class. Egyptian Cat vs. Persian Cat: predicted label is 283, corresponding to "Persian cat"; target label is 285, corresponding to "Egyptian cat". We adopt this setting to perform a transition from the “Persian cat” class to the “Egyptian cat” class. For a detailed mapping between classes and their corresponding labels, please refer to this link.
Hyperparameters Settings We provide the detailed hyperparameter settings for each dataset in Table 6. $\alpha$ is the weight of the adversarial defense loss, $\lambda$ is the weight of the $l_{1}$ loss used during mask generation, and $\gamma$ is the threshold for generating the binary mask.
@@ A. Training Details
CelebA-HQ Smile 0.05 0.001 0.1 50 500
CelebA-HQ Age 0.05 0.001 0.1 50 500
@@
Sorrel-Zebra 0.1 0.001 0.15 25 1000
Cougar-Cheetah 0.1 0.001 0.15 25 1000
Persian Cat-Egyptian Cat 0.05 0.001 0.15 25 1000
B. Additional Experimental Results

### v2

@@ 4.2 Causally Post Refinement
[…]
@@
The mask indicates which pixels need to be modified through $l_{1}$ loss and which do not. Intuitively, we need to combine $x^{\prime}$ generated in the last subsection and the original image $x$ with the mask to achieve causal perturbation in CEs with well-preserved pixels in the rest parts. However, directly combining $x^{\prime}$ and $x$ using the mask is not feasible, as it would result in clearly visible region boundaries in the CE. In our work, we leverage the image inpainting technique (Lugmayr et al., 2022) to tackle this problem. The inpainting process is shown as follows:
@@ 5.1 Experimental Setup
[…]
Baselines We selected four recent and state-of-the-art methods as our baselines: STEEX (Jacob et al., 2022), DiVE (Rodriguez et al., 2021), DiME (Jeanneret et al., 2022), and ACE (Jeanneret et al., 2023). Among them, DiVE includes an extended variant called DiVE100, in which the number of optimization steps is reduced to 100.
Models Weights and Architectures To ensure a fair comparison, we use the same pre-trained DDPM weights as those used in the baselines for each dataset. For ImageNet, we use the ResNet-50 model pre-trained by PyTorch (Paszke, 2019) as the classifier. For the other datasets, we use DenseNet-121 as the classifier and adopt the pre-trained weights provided by the baselines. Additional training details, including hyperparameter settings, are provided in Appendix A.
Evaluation Metrics We evaluate our method and the baselines across multiple dimensions, each assessed using several metrics. The first evaluation dimension is validity, which we measure using Flip Rate (FR) and Counterfactual Transition (COUT). FR denotes the proportion of samples for which the counterfactual successfully causes the classifier to change its prediction. COUT measures the transition scores between the original and counterfactual images, with values ranging from 1-1 to $1$. The second evaluation dimension focuses on sparsity and proximity, which we assess using Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Correlation Difference (CD) (Jeanneret et al., 2022), and SimSiam Similarity (S3) (Chen & He, 2021). LPIPS measures the perceptual similarity between two images based on human visual perception. CD computes the average number of non-target attributes that are unintentionally altered across all samples. S3 evaluates the cosine similarity between the original and counterfactual images using SimSiam as the encoding network. The third evaluation dimension is realism, which we measure using Fréchet Inception Distance (FID) (Heusel et al., 2017) and split FID (sFID) (Jeanneret et al., 2023). Since counterfactual images typically share most pixels with the original images, FID tends to be underestimated and may fail to reflect their true differences. To address this issue, sFID is proposed. It splits both the original and counterfactual image sets into two subsets, computes FID in a cross-wise manner, and repeats this process ten times to obtain an average score. All results reported in this paper are averaged over multiple experimental runs. We omit standard deviations as they are negligible compared to the corresponding mean values.
Table 1: Evaluation Results on CelebA-HQ: Smile.
@@
DiVE - - - - - 107.0 -
STEEX - - - - - 21.9 -
DiME 90.00% -0.2797 0.1689 0.9245 2.52 15.8 26.79
ACE 77.50% 0.2369 0.0353 0.9868 2.76 5.9 38.20
Ours 97.85% 0.6751 0.0331 0.9889 2.42 3.9 26.58
Table 2: Evaluation Results on CelebA-HQ: Age
@@
DiVE - - - - - 107.5 -
STEEX - - - - - 26.8 -
DiME 94.55% -0.3339 0.1689 0.9202 4.00 17.8 26.54
ACE 63.45% 0.0567 0.0448 0.9876 5.73 9.7 33.95
Ours 96.15% 0.5489 0.0466 0.9893 5.33 5.6 27.84
@@
Datasets Methods FR $\uparrow$ COUT $\uparrow$ LPIPS $\downarrow$ S3 $\uparrow$ FID $\downarrow$ sFID $\downarrow$
Sorrel-
Zebra
ACE 76.67% 0.0967 0.1670 0.9285 77.76 124.66
Ours 93.33% 0.2219 0.0468 0.9338 72.26 116.28
Cougar-
Cheetah
ACE 76.67% 0.2059 0.1809 0.9625 108.67 147.79
Ours 100.00% 0.4186 0.0470 0.9616 102.56 137.91
Persian Cat-
Egyptian Cat
ACE 80.00% 0.2326 0.1531 0.8939 139.65 227.07
Ours 83.33% 0.2660 0.0466 0.8744 134.89 215.07
@@ 5.2 Quantitative Analysis
In this section, we conduct a quantitative analysis of all datasets with results summarized in Table 1 through Table 6. We highlight the best performance in bold and the second-best in underline. The results for the baselines DiVE, DiVE100, and STEEX, we extract the results from the original papers. For the data not reported in the original papers, we leave the corresponding entries blank in the tables.
Tables 1and  2 show the results on the CelebA-HQ Smile and Age dataset, Table 3 shows the results on the ImageNet dataset. As observed, our method achieves the best performance on the vast majority of metrics. More specifically, our method achieves the highest FR and COUT scores, while also attaining the lower LPIPS and higher S3 values, along with relatively low FID and sFID scores. This indicates that our method can achieve the highest flip rate with minimal modifications. In other words, it attains the highest validity while preserving both sparsity and proximity. Although in Table 2 our method performs slightly worse than DiME on certain metrics, the COUT of DiME is negative. This indicates that in DiME’s method, the classifier lacks sufficient confidence in label flipping. Even though the other metrics may show marginal improvements, they do not necessarily indicate that the counterfactual images are of higher quality than those produced by our method. Due to page limitations, we present the results on the other datasets in Appendix B.
Figure 3: Qualitative results on CelebA (Smile & Age). Red circles highlight unintended changes produced by ACE, while green circles indicate the improvements made by our method.
@@ 5.3 Qualitative Analysis
In this section, we perform case studies to straightforwardly evaluate the quality of the generated counterfactual images from a human perceptual perspective. For CelebA datasets, we compare the counterfactual images generated by our approach with those produced by the best-performing method, ACE, to demonstrate the visual superiority of our method. As shown in Figure 3, ACE introduces inexplicable color and shape distortions, as well as unintended attributes in the generated images. For example, in the transformation from “young” to “old,” some female samples are incorrectly rendered with dark patches on the chin—an artifact caused by the spurious correlation between beard and age in male samples. This results in unmeaningful CEs. In contrast, our method effectively mitigates such artifacts by avoiding reliance on spurious cues. Figure 4 shows the results on ImageNet. As our method and ACE yield visually indistinguishable outputs on this dataset, we restrict the comparison to the original images and their counterfactual counterparts. Qualitative results for other datasets are provided in Appendix B.
@@ 5.4 VLM Evaluations
Vision-Language Models (VLMs), through large-scale pretraining on image–text pairs, demonstrate strong reasoning capabilities in tasks such as visual question answering and image captioning (Bordes et al., 2024). Overcoming the limitation of traditional CE metrics that only provide shallow numerical evaluations and fail to capture the semantic meaning of images, VLMs provide more human-like evaluations by leveraging their rich prior knowledge in pertaining (Huang et al., 2025). To better suit the specific nature of counterfactual image generation, we carefully design unique prompts with evaluation criteria covering validity, sparsity, proximity, and realism. The prompt design for each dataset can be found in Appendix C. Notably, sparsity in our prompt is partially assessed from a causal perspective, examining whether the modified regions in the image are necessary or causally relevant, while ensuring that unrelated parts remain unchanged. In the context of our task, we utilize GPT-4o to assign scores ranging from 1 to 10 for each of the four evaluation dimensions. Given that these dimensions hold equal importance in counterfactual generation, the final score is computed as the average across all four. As shown in Table 4, our method outperforms the baselines in 7 out of the 8 sub-datasets. Combined with the powerful visual and textual reasoning capabilities of VLMs, we have strong reason to believe that the counterfactual images generated by our method offer the highest level of interpretability.
@@ 5.5 Parameters Study
Figure 4: Line plots of FR and COUT with respect to varying values of $\alpha$.
In this section, we analyze the impact of the hyperparameter $\alpha$ on counterfactual image generation. We present results on the CelebA: Smile dataset, while results on the other datasets are provided in Appendix D. As shown in Figure 4, when $\alpha$ increases, the model is progressively constrained from modifying spurious factors. This trend leads to a monotonic decrease in both the flip rate and the COUT score. The reduction in flip rate indicates that many successful counterfactual transitions at lower $\alpha$ values are largely driven by spurious shortcuts; once these features are restricted, generating effective counterfactuals becomes more challenging. The concurrent decline in COUT further suggests that the transition trajectories become less smooth, with the classifier maintaining high confidence in the original class and showing limited confidence shifts toward the target class throughout the interpolation path. Taken together, these results demonstrate that higher $\alpha$ enforces a stricter focus on causal features, thereby producing counterfactuals that are harder to realize but more faithful to causal semantics. Therefore, it is crucial to strike a balance between flip effectiveness and visual realism. Fortunately, with a properly chosen $\alpha$, our method effectively achieves this balance, producing counterfactuals with both high flip rates and high image quality.
@@ 6 Conclusion
In this paper, we propose a novel framework, CECAS, which innovatively designs a causally-guided adversarial method to generate counterfactual visual explanations that constrain less informative changes on spurious factors. The framework can be applied to any DNN models. Experimental results demonstrate that our method consistently outperforms existing approaches in quantitative evaluation, qualitative analysis, and VLM-based assessment. CECAS achieves a better balance among the four key properties of counterfactual explanations: validity, sparsity, proximity, and realism. Therefore, we argue that the counterfactual explanations generated by our method offer a more causally grounded understanding of the classifier’s decision-making and prediction mechanisms.
@@
Appendix A Training Details
A.1 Data Preparation
The ten datasets discussed in the main text are categorized into two types: binary datasets and multi-label datasets. CelebA Smile, CelebA Age, CelebA Gender, CelebA-HQ Smile, CelebA-HQ Age, CelebA-HQ Gender, and BDD-OIA belong to the first category, while Cougar vs. Cheetah, Sorrel vs. Zebra, and Egyptian Cat vs. Persian Cat fall into the latter. For binary-label datasets, if the original predicted label is $y$ ($y$ can either be 0 or 1), then the target label is set to 1y1-y. For multi-label datasets, since ImageNet contains 1,000 classes, the prediction and target labels differ for each dataset subset. Specifically, Cougar vs. Cheetah: predicted label is 286, corresponding to “cougar”; target label is 293, corresponding to ”cheetah”. We adopt this setting to perform a transition from the “cougar” class to the “cheetah” class. Sorrel vs. Zebra: The predicted label is 339, which corresponds to ”sorrel”; the target label is 340, which corresponds to ”zebra”. We adopt this setting to perform a transition from the “sorrel” class to the “zebra” class. Egyptian Cat vs. Persian Cat: predicted label is 283, corresponding to ”Persian cat”; target label is 285, corresponding to ”Egyptian cat”. We adopt this setting to perform a transition from the “Persian cat” class to the “Egyptian cat” class.
A.2 Hyperparameters Settings
We provide the detailed hyperparameter settings for each dataset in Table 5. $\alpha$ is the weight of the adversarial defense loss, $\lambda$ is the weight of the $l_{1}$ loss used during mask generation, and $\gamma$ is the threshold for generating the binary mask.
@@ A.2 Hyperparameters Settings
CelebA Gender 0.4 0.001 0.1 50 500
CelebA-HQ Smile 0.01 0.001 0.1 50 500
CelebA-HQ Age 0.03 0.001 0.1 50 500
CelebA-HQ Gender 0.005 0.001 0.1 50 500
@@
Sorrel-Zebra 0.03 0.001 0.15 100 1000
Cougar-Cheetah 0.02 0.001 0.15 100 1000
Persian Cat-Egyptian Cat 0.04 0.001 0.15 100 1000
Appendix B Additional Experimental Results